Can ∫ (x^6 +8)^2 dx be integrated with u = x^6+8? Explain

1
Expert's answer
2022-05-16T10:03:07-0400

"u=x^6+8" is a differentiable function. Its rang is "(-\infin, \infin)."

The fuction"f(x)=(x^6+8)^2" is continuous on "(-\infin, \infin)."

Then we can use the Substitution Rule and


"u=x^6+8, du=6x^5 dx"

"\int(x^6+8)^2 (6x^5)dx=\int u^2du=\dfrac{u^3}{3}+C"

"=\dfrac{(x^6+8)^3}{3}+C"

In our case we have

"\int(x^6+8)^2 dx"

instead of


"\int(x^6+8)^2 (6x^5)dx"

Therefore it is not useful to integrate

"\int(x^6+8)^2 dx"

with "u-" substitution "u=x^6+8."


"\int(x^6+8)^2 dx=\int(x^{12}+16x^6+64) dx"

"=\dfrac{x^{13}}{13}+\dfrac{16x^7}{7}+64x+C"